b3lla2025:

In the equation 5y = 3x - 15 if I plug 0(zero) in for x what would y be? (answer is just the number like 15) *

supie:

So if the plug in 0 for x we would have 5y=(3)(0)−15 Yes?

b3lla2025:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie So if the plug in 0 for x we would have 5y=(3)(0)−15 Yes? \(\color{#0cbb34}{\text{End of Quote}}\) yes I think so

supie:

So do you know how we would solve that?

b3lla2025:

would it be - 12

supie:

y=-12?

b3lla2025:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie y=-12? \(\color{#0cbb34}{\text{End of Quote}}\) maybe I'm confused

supie:

hm well the first thing we would to is multiply 3 and 0 which equals 0 So we would have \(5y=0+−15\) thne we combine like terms you think you would do that?

supie:

then*

supie:

\([5y=(0+−15)]\) Solve what it inside of the parenthesis `()`

b3lla2025:

5y- 15

supie:

Yes, do you know the next step?

b3lla2025:

then add 0

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @b3lla2025 5y- 15 \(\color{#0cbb34}{\text{End of Quote}}\) Actually it's \(5y=−15\)

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @b3lla2025 then add 0 \(\color{#0cbb34}{\text{End of Quote}}\) No-

supie:

What we do is ÷5 to both sides

b3lla2025:

so -15 divided by 5 =-3

b3lla2025:

so y=-3

supie:

So that should look like \(\LARGE \frac{5y}{5}=\frac{-15}{5}\) So then 5y/5 cancels \(\cancel{\LARGE \frac{5y}{5}}\LARGE=\frac{-15}{5}\) So all you have left is \(\frac{-15}{5}\) so we simplify that then we get \(\frac{-15}{5}=-3\) So y=-3

b3lla2025:

so I was right